OpenStudy (anonymous):

m, 2m, 4m, . . . The first term in the sequence above is m, and each term thereafter is equal to twice the previous term. If m is an integer, which of the following could NOT be the sum of the first four terms of this sequence?

OpenStudy (anonymous):

a) -25 b)-15 c)45 d) 75 e)120

OpenStudy (anonymous):

The first four terms are m, 2m, 4m and 8m Plug in easy numbers for "m" to see what the sum is. Forexample, if m= 1, the sum is 1 + 2 + 4 + 8 = 15. If m= 2, the sum is 2 + 4 + 8 + 16 = 30. You will see that the sums are always multiples of 15, so answer A is correct since it isn’t a multiple of 15. so, m+ 2,m+ 4,m+ 8,m= 15m so that the sum of the first four terms is always a multiple of 15.

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

yes ! thank you so much